Gonzalez, a U.S. citizen, owns 1% of USAco, a domestic corporation. All other shareholders of USAco are unrelated foreign persons. In a Type B reorganization, FORco transfers shares of its voting stock to USAco shareholders in exchange for 100% of the stock in USAco. Gonzalez realizes a gain on the exchange. As a renowned and reputable international tax planning consultants, Gonzalez retains your services so that he can minimize any potential tax liabilities.

Relying on the concepts from TCJA:

  • How he can avoid the outbound-toll charge on the gain?.
Answer #1

100,00,000*35/100= 35,00,000 pre credit tax

10,00,000*50/100= 5,00,000 foreign tax paid

Total US tax OWNED= 35,00,000-5,00,000= 30,00,000

ETR = 30,00,000/1,00,00,000= 30%

in this way he can avoid the outbound toll charge on the gain